Please help solve :
fraction:9/2 -n/12=1/4

Answer by checkley71(8403)

9/2-n/12=1/4
-n/12=1/4-9/2
-n/12=(1-2*9)/4
-n/12=(1-18)/4
-n/12=-17/4 now cross multiply
-4n=-17*12
-4n=-204
n=-204/-4
n=51 answer.